# Resonant amplification of multi-messenger emission in rotating stellar core collapse

### v1

@@ Abstract
In a series of axisymmetric core-collapse supernova simulations extending up to $\sim 2\,\mathrm{s}$, we identify a regime of pre-collapse central rotation rates ($\sim 1\,\mathrm{Hz}$) that greatly enhances the emission of gravitational waves (GWs) during extended periods of time after bounce. The enhancement is a consequence of the resonance between the frequency of the fundamental quadrupolar ${}^{2}f$-mode of oscillation of the proto-neutron star and the frequency of the epicyclic oscillations at the boundary of the inner core. We observe periods of about several hundred milliseconds each where the resonance is active. The GW emission enhancement produces a correlated resonant modulation of the associated neutrino signal at the same frequencies. With GW frequencies of $\mathcal{O}(1\,\mathrm{kHz})$ and strain amplitudes within the sensitivity curves of current and next-generation interferometers at distances of $\mathcal{O}(1\,\mathrm{Mpc})$, this resonant-amplification mechanism may represent a potential game-changer for unveiling the supernova explosion mechanism through multi-messenger astronomy.
@@ I Introduction
Core collapse supernovae (CCSNe) mark the final stage in the evolution of massive stars ($M_{\textnormal{ZAMS}}\gtrsim 8\,\mathrm{M_{\odot}}$) and are among the prime sources of multimessenger emission as they release gravitational waves (GWs), neutrinos, and electromagnetic signals. The detection of 25 neutrinos from Supernova 1987A [1, 2], along with its electromagnetic counterpart [3, 4], demonstrated that multimessenger observations of CCSNe occurring in the Milky Way are feasible.
@@
The rotation of the progenitor star is pivotal in shaping the waveform of the resulting GW signal and, consequently, the detectability of the event [28, 29]. Specifically, when the core of the progenitor star achieves a certain rotational speed, resonances can develop between the PNS oscillation modes and the core’s rotational frequency. Such resonances have been identified in simplified 2-dimensional (2D) models of rotating relativistic stars [30], in the collapse of neutron stars to strange quark stars triggered by phase transitions [31], and during the early phases ($\lesssim 100\,\mathrm{ms}$ post-bounce) of CCSNe simulations [32]. However, the full evolution of these resonances through the entire operation of the CCSNe engine remains to be comprehensively explored. Previous studies indicate that resonances can notably amplify the GW amplitude, introduce new oscillatory modes, and imprint corresponding signatures in the neutrino signal [32].
In this Letter we delve into the rich tapestry of multimessenger signals emerging from CCSN simulations with different initial rotation rates over a span of $\sim 2\,$s. Our findings reveal that when the stellar core rotates at an intermediate rate ($\approx 1\,\mathrm{Hz}$), it triggers resonant modes in the PNS that imprint synchronized modulations on the neutrino signal. Remarkably, these resonant GWs generate colossal strains, even surpassing those at bounce, and manifest within mere hundreds of milliseconds after collapse. This striking discovery not only deepens our understanding of the dynamic interplay in stellar explosions but also paves the way for novel insights into their multimessenger signatures.
@@ II Methods
We performed 2D CCSN simulations using the Aenus-ALCAR code [33, 34, 35] that couples special relativistic magnetohydrodynamics with a spectral two-moment neutrino transport scheme and incorporates an approximately relativistic gravitational potential [36]. For densities exceeding $8\times 10^{7}\,\mathrm{g/cm^{3}}$, we employed the SFHo EOS [37], accounting for photons, electrons, positrons, nucleons, and a full ensemble of light and heavy nuclei in nuclear statistical equilibrium. The GW signal was extracted using the quadrupole formula [38]. The simulations were performed in spherical coordinates on grids with 480 logarithmically spaced radial zones and 128 linearly spaced angular zones, while neutrino energies spanning from $0\,\mathrm{MeV}$ to $440\,\mathrm{MeV}$ were resolved with 12 logarithmically spaced bins. Each run was evolved for $2\,\mathrm{s}$, commencing from the pre-collapse phase.
After an extensive survey of rotating stellar models evolved to several hundreds of milliseconds post-collapse, we identified a progenitor with a particularly pronounced resonant coupling between GWs and rotation. Specifically, it is a Wolf-Rayet star of subsolar metallicity ($Z=\mbox{0.02}\,\mathrm{Z_{\odot}}$) and zero-age main sequence mass of $M_{\textnormal{ZAMS}}=17\,\mathrm{M_{\odot}}$, evolved with both rotation and magnetic fields up to the pre-collapse stage [39]. Its initially high rotational velocity, coupled with enhanced rotational mixing, drives chemically homogeneous evolution. Building on this model, we explored three rotational configurations: the original (SR), with central rotational rate of $\Omega_{\rm c}=0.29\,\mathrm{rad/s}$, an intermediate case (IR) with a rate amplified by a factor of 3.5, and a fast-rotating scenario (FR) with a 12-fold increase. Although these increased rotation rates are substantial, they remain well within the computed range for massive stars [40, 41].
@@ III Results
The plus polarisation of the GW signal, $h_{+}$, for a source at distance $\mathcal{D}$ (see top panel of Figure 1) reveals the brief yet intense bounce signal characteristic of rapidly rotating cores [43, 44]. In the most rapidly rotating model FR, the signal amplitude reaches $\mathcal{D}h_{+}\gtrsim 100\,\mathrm{cm}$. Following the bounce, the waveform transitions into a regime of weaker oscillations that remain centered around zero in model SR or gradually drift towards positive values in models IR and FR, which ultimately explode via polar jets.
@@
The bottom panel of Figure 1 shows the GW spectrogram of model IR, computed using a short-time Fourier transform with a $10\,\mathrm{ms}$ time window. The two GW bursts display distinct spectral evolutions: the first rises from $\sim 550\,\mathrm{Hz}$ to $\sim 1\,\mathrm{kHz}$, while the second steadily hovers near $1\,\mathrm{kHz}$. These frequencies align with the predicted fundamental quadrupolar ${}^{2}f$-mode (green line; see, e.g., [42, 27]). Moreover, additional harmonics appear at twice and thrice these frequencies, albeit with progressively lower amplitudes.
@@
Using the quadrupole formula, we pinpoint the origin of the GW strain to the innermost part of the star (hereafter PNS core), which we defined as the region within the iso-entropy surface at $4\,\mathrm{k_{B}/bry}$ [45], corresponding to a radius of $r_{\textnormal{core}}\sim 20\,\mathrm{km}$). The pronounced high-amplitude emission arises from a resonant interplay between PNS modes and rotation. The centrifugal force causes fluid elements within the PNS core to oscillate in the direction of the cylindrical radial distance. To quantify these oscillations, we use the epicyclic frequency, defined as
@@
Within the PNS core, we focus on a conical region spanning the colatitude range $[15\,\mathrm{{}^{\circ}},85\,\mathrm{{}^{\circ}}]$. At each radius, we determine the maximum epicyclic frequency and subsequently compute the average and standard deviation of these maxima. This specific colatitude range was chosen to exclude the polar region (prone to numerical artifacts in axial symmetry) and the equator, where converging convective flows compromise the accuracy of the ${}^{n}f_{\rm epi}$ evaluation. Recognizing that simply taking the maximum may lead to an overestimate, we conservatively define the lower uncertainty as the largest value between the standard deviation and 10% of the maximum of the epicyclic frequency. The bottom panel of Figure 1 shows the evolution of these averages (blue lines) and uncertainties (shades) for ${}^{1}f_{\textnormal{epi}}$ (solid), ${}^{2}f_{\textnormal{epi}}$ (dashed), and ${}^{3}f_{\textnormal{epi}}$ (dotted).
@@
Resonance occurs when the fundamental epicyclic frequency intersects or comes sufficiently close to the ${}^{2}f$-mode frequency. To give a first analytic estimate of the latter we employ the quasi-universal relations of [42]. Although these relations were obtained from non-rotating 1D CCSN simulations and thus may exhibit minor deviations from our models, they still provide a useful point of reference for comparison with the epicyclic frequency.
@@
To further validate the resonance mechanism, we resort to the computed data to give a second estimate of the frequency of the quasi-radial ($l=0$) and quadrupolar ($l=2$) modes. As in [30], we use as proxies for the amplitudes of the $l=0$ and $l=2$ modes the real parts of the Fourier spectra of the mass density at $r=25\,\mathrm{km},\,\theta=\pi/2$ and of $v_{\theta}$ at $r=25\,\mathrm{km},\theta=\pi/4$, respectively. In Figure 2 we present the Fourier transform amplitude of the GW, density and $v_{\theta}$, each normalized by their maximum during the second resonant phase ($1.15-1.45\,\mathrm{s}$). The GW spectrum (green line) exhibits a series of distinct peaks starting at $\sim 1\,\mathrm{kHz}$, spaced at intervals of $\sim 1\,\mathrm{kHz}$. These peaks persist up to $\sim 7\,\mathrm{kHz}$, though their amplitude steadily diminishes with increasing frequency, a behaviour indicative of overtones of the epicyclic frequency.
@@
Figure 3: Characteristic GW spectra for models SR (red), IR (black), and FR (yellow), assuming a source distance of $1\,\mathrm{Mpc}$, compared with the design sensitivity curves of current and next-generation interferometers.
@@
$h_{\textnormal{char}}(f)=\frac{1}{\pi\mathcal{D}}\sqrt{\frac{2G}{c^{3}}\frac{\textnormal{d}E}{\textnormal{d}f}},$ (2)
where $c$ is the speed of light, $G$ is the gravitational constant, and $\frac{\textnormal{d}E}{\textnormal{d}f}$ denotes the time-integrated energy spectrum. In Figure 3 we show $h_{\textnormal{char}}/f^{1/2}$ for models SR, IR, and FR at $\mathcal{D}=1\,\mathrm{Mpc}$, juxtaposed with the amplitude spectral density (ASD) of current and next-generation interferometers, namely Advanced LIGO [47, 48], Advanced Virgo [47, 48], KAGRA [47, 48], ET [49, 50], and CE [51, 52].
@@
Model IR differs significantly from the others. It displays two low-frequency spectral humps: one at intermediate frequencies ($60-200\,\mathrm{Hz}$, and another below $30\,\mathrm{Hz}$. The most striking feature of IR is the unmistakable resonance imprint on its characteristic strain. This model yields the strongest signal, with a peak strain more than twice that observed in SR–translating into a peak ASD more than 10 (100) times larger than in model SR (FR)-(see Figure 3). Four prominent peaks, spanning from $\sim 600\,\mathrm{Hz}$ to $\sim 1\,\mathrm{kHz}$, mark the resonant frequencies during the first and second resonant phases (see Figure 1). Remarkably, even at a distance of $1\,\mathrm{Mpc}$, these features lie within the detection range of current interferometers like Advanced LIGO and Virgo, with next-generation detectors promising even greater sensitivity. Overtones at $f\gtrsim 2\,\mathrm{kHz}$ and $f\gtrsim 3\,\mathrm{kHz}$ are also present.
@@
[…]

### v2

# Resonant amplification of multimessenger emission in rotating stellar core collapse

@@ Abstract
In a series of axisymmetric core-collapse supernova simulations extending up to $\sim 2\,\mathrm{s}$, we identify a regime of pre-collapse central rotation rates ($\sim 1\,\mathrm{Hz}$) that greatly enhances the emission of gravitational waves (GWs) during extended periods of time after bounce. The enhancement is a consequence of the resonance between the frequency of the fundamental quadrupolar $^{2}f-$mode of oscillation of the proto-neutron star and the frequency of the epicyclic oscillations at the boundary of the inner core. We observe periods of about several hundred milliseconds each where the resonance is active. The GW emission enhancement produces a correlated resonant modulation of the associated neutrino signal at the same frequencies. With GW frequencies of $\mathcal{O}(1\,\mathrm{kHz})$ and strain amplitudes within the sensitivity curves of current and next-generation interferometers at distances of $\mathcal{O}(1\,\mathrm{Mpc})$, this resonant-amplification mechanism may represent a potential game-changer for unveiling the supernova explosion mechanism through multimessenger astronomy.
@@ I Introduction
Core collapse supernovae (CCSNe) mark the final stage in the evolution of massive stars ($M_{\textnormal{ZAMS}}\gtrsim 8\,\mathrm{M_{\odot}}$) and are among the prime sources of multimessenger emission as they release gravitational waves (GWs), neutrinos, and electromagnetic signals. The detection of 25 neutrinos from Supernova 1987A [1, 2, 3], along with its electromagnetic counterpart [4, 5], demonstrated that multimessenger observations of CCSNe are feasible.
@@
The rotation of the progenitor star is pivotal in shaping the waveform of the resulting GW signal and, consequently, the detectability of the event [29, 30]. Specifically, when the core of the progenitor star achieves a certain rotational speed, resonances can develop between the PNS oscillation modes and the core’s rotational frequency. Such resonances have been identified in simplified 2-dimensional (2D) models of rotating relativistic stars [31], and in the collapse of neutron stars to strange quark stars triggered by phase transitions [32]. Additionally, during the early phases ($\lesssim 100\,\mathrm{ms}$ post-bounce) of CCSNe simulations, rotation can excite specific PNS modes [33]. However, the full evolution of these resonances through the entire operation of the CCSNe engine remains to be comprehensively explored. Previous studies indicate that particular pre-collapse rotation rates can notably amplify the GW amplitude, introduce new oscillatory modes, and imprint corresponding signatures in the neutrino signal [33].
In this work we delve into the rich tapestry of multimessenger signals emerging from CCSN simulations with different initial rotation rates over a span of $\sim 2\,$s. Our findings reveal that when the stellar core rotates at an intermediate rate ($\approx 1\,\mathrm{Hz}$), it triggers resonant modes in the PNS that imprint synchronized modulations on the neutrino signal. Remarkably, these resonant GWs generate substantial strains, even surpassing those at bounce, and manifest within mere hundreds of milliseconds after collapse. This striking discovery not only deepens our understanding of the dynamic interplay in stellar explosions but also paves the way for novel insights into their multimessenger signatures.
The remainder of this paper is organised as follows: Section II outlines our numerical setup and presents the progenitor model, initial rotation profiles, and nuclear EOS used to perform the CCSN simulations. In Section III we present the main findings of our study. Finally, in Section IV we discuss our results and draw our conclusions.
@@ II Methods
We performed 2D CCSN simulations using the Aenus-ALCAR code [34, 35, 36] that couples special relativistic magnetohydrodynamics with a spectral two-moment neutrino transport scheme and incorporates an approximately relativistic gravitational potential [37]. For densities exceeding $8\times 10^{7}\,\mathrm{g/cm^{3}}$, we employed the SFHo EOS [38], accounting for photons, electrons, positrons, nucleons, and a full ensemble of light and heavy nuclei in nuclear statistical equilibrium. The GW signal was extracted using the quadrupole formula [39]. The simulations were performed in spherical coordinates on grids with 480 logarithmically spaced radial zones and 128 linearly spaced angular zones, while neutrino energies spanning from $0\,\mathrm{MeV}$ to $440\,\mathrm{MeV}$ were resolved with 12 logarithmically spaced bins. Each run was evolved for $2\,\mathrm{s}$, commencing from the pre-collapse phase. In the Appendix A we assess the dependence of our results with increased grid resolution.
After an extensive survey of rotating stellar models evolved to several hundreds of milliseconds post-collapse, we identified several simulations that presented resonant coupling between GWs and rotation. From these, we selected one case in which this feature is particularly pronounced. Specifically, it is a Wolf-Rayet star of subsolar metallicity ($Z=\mbox{0.02}\,\mathrm{Z_{\odot}}$) and zero-age main sequence mass of $M_{\textnormal{ZAMS}}=17\,\mathrm{M_{\odot}}$, evolved with both rotation and magnetic fields up to the pre-collapse stage [40]. Its initially high rotational velocity, coupled with enhanced rotational mixing, drives chemically homogeneous evolution. Building on this model, and reconstructing the magnetic field topology from the saturation strengths originally included in the (one-dimensional) stellar evolution model (of the order of $10^{9}\,$G inside the iron core), we explored three rotational configurations: the original (SR), with central rotational rate of $\Omega_{\rm c}=0.29\,\mathrm{rad/s}$, an intermediate case (IR) with a rate amplified by a factor of 3.5, and a fast-rotating scenario (FR) with a 12-fold increase. Although these increased rotation rates are substantial, they remain well within the computed range for massive stars [41, 42].
@@ III Results
The plus polarisation of the GW signal, $h_{+}$, for a source at distance $\mathcal{D}$ (see top panel of Figure 1) reveals the brief yet intense bounce signal characteristic of rapidly rotating cores [44, 45]. In the most rapidly rotating model FR, the signal amplitude reaches $\mathcal{D}h_{+}\gtrsim 100\,\mathrm{cm}$. Following the bounce, the waveform transitions into a regime of weaker oscillations that remain centred around zero in model SR or gradually drift towards positive values in models IR and FR, which ultimately explode via polar jets.
@@
The bottom panel of Figure 1 shows the GW spectrogram of model IR, computed using a short-time Fourier transform with a $10\,\mathrm{ms}$ time window. The two GW bursts display distinct spectral evolutions: the first rises from $\sim 550\,\mathrm{Hz}$ to $\sim 1\,\mathrm{kHz}$, while the second steadily hovers near $1\,\mathrm{kHz}$. The frequencies and time evolution of the most intense ascending spectral feature align with the predicted fundamental quadrupolar $^{2}f-$mode (green line computed using as independent variable $x=M_{\rm pns}/R_{\rm pns}^{3}$ instead of the $x=M_{\rm shock}/R_{\rm shock}^{3}$ in the fits provided by [26]). Moreover, additional power in the spectrogram growing in parallel to the main mode and resembling higher harmonics of the former one appear at twice and thrice these frequencies, albeit with progressively lower amplitudes. Both facts strongly suggest that the rising modes in the spectrogram likely result from PNS $f$-modes. We stress the fact that the green line in Figure 1 must be regarded as a proxy of the actual $^{2}f-$mode frequency. Hence, we refer to this proxy as approximate $^{2}f$ mode.
Figure 2: Snapshots of the radial velocity illustrating a full oscillation cycle of the PNS at (clockwise from the top-left panel) $1283\,\mathrm{s}, 1.1$$285\,\mathrm{s}, 1.1$$1287\,\mathrm{s}$, and $1.1291\,\mathrm{s}$. Each of the panels shows one hemisphere only. Black lines represent the PNS radius, defined as the isodensity line at $10^{11}\,\mathrm{g/cm^{3}}$. Streamlines represent the velocity field. Background colours indicate inward (blue) and outward-moving (red) matter.
@@
Using the quadrupole formula including surface terms (see, e.g., [46]), we pinpoint the origin of the GW strain to the innermost part of the star (hereafter PNS core), which we defined as the region within the iso-entropy surface at $4\,\mathrm{k_{B}/baryon}$ [47], corresponding to a radius of $r_{\textnormal{core}}\sim 20\,\mathrm{km}$). The pronounced high-amplitude emission arises from a resonant interplay between PNS modes and rotation. The centrifugal force causes fluid elements within the PNS core to oscillate in the direction of the cylindrical radial distance. To quantify these oscillations, we use the epicyclic frequency, defined as
@@
Within the PNS core, we focus on a conical region spanning the colatitude range $15\,\mathrm{{}^{\circ}},85\,\mathrm{{}^{\circ}}]$. At each radius, we determine the maximum epicyclic frequency and subsequently compute the average and standard deviation of these maxima. This specific colatitude range was chosen to exclude the polar region (prone to numerical artifacts in axial symmetry) and the equator, where converging convective flows compromise the accuracy of the $^{n}f_{\rm epi}$ evaluation. Recognizing that simply taking the maximum may lead to an overestimate, we conservatively define the lower uncertainty as the largest value between the standard deviation and 10% of the maximum of the epicyclic frequency. The bottom panel of Figure 1 shows the evolution of these averages (blue lines) and uncertainties (shades) for $^{1}f_{\textnormal{epi}}$ (solid), $^{2}f_{\textnormal{epi}}$ (dashed), and $^{3}f_{\textnormal{epi}}$ (dotted).
@@
Resonance occurs when the fundamental epicyclic frequency intersects or comes sufficiently close to the $^{2}f-$mode frequency. To give a first analytic estimate of the latter we employ the quasi-universal relations of [26]. These relations were obtained from non-rotating 1D CCSN simulations. Thus, their approximate formulae may exhibit deviations from our models, yet they provide a rough reference for comparison with the epicyclic frequency.
During the resonance phases, the PNS core oscillates violently over few tens of milliseconds. Since the PNS core and PNS essentially coincide during the second resonant phase, the two terms are used interchangeably in the following discussion. In Figure 2 we provide a time-sequence visualization of the fluid motion within the PNS core and the associated deformation during the second resonant phase, showing the radial velocity ($v_{r}$) during a full oscillation cycle. The PNS core motion takes the form of a large-scale oscillation in which the core undergoes over a full cycle the following dynamical sequence: a roughly isotropic contraction (top left), formation of two vortices, one per hemisphere, with inward/outward radial velocity at the equator/poles (top right), expansion (bottom right), and oppositely oriented vortex pair (bottom left). Over the course of one oscillation, the PNS core changes its quadrupole moment due to the differential between motion in the equatorial and polar regions (second and fourth phase). At the outer border of the PNS (black line), each cycle of the oscillation launches a wave into the post-shock region, visible in the four panels as the approximately isotropic patterns of alternating positive and negative velocities moving away from the PNS.
@@
To further validate the resonance mechanism, we resort to the computed data to give a second estimate of the frequency of the quasi-radial ($l=0$) and quadrupolar ($l=2$) modes. As in [31], we use as proxies for the amplitudes of the $l=0$ and $l=2$ modes the real parts of the Fourier spectra of the mass density at $r=25\,\mathrm{km},\,\theta=\pi/2$ and of $v_{\theta}$ at $r=25\,\mathrm{km},\theta=\pi/4$, respectively. The choice of $r=25\,\mathrm{km}$ ensures that the extracted signal reflects fluid displacements induced by the oscillations of the PNS core. In Figure 3 we present the Fourier transform amplitude (FTA) of the density and $v_{\theta}$, each normalized by their maximum during the second resonant phase ($1.15-1.45\,\mathrm{s}$), along with the GW FTA computed in the whole domain ($r\leq 10^{10}\,$cm). The GW spectrum (green line) exhibits a series of distinct peaks starting at $\sim 1\,\mathrm{kHz}$, spaced at intervals of $\sim 1\,\mathrm{kHz}$. These peaks persist up to $\sim 7\,\mathrm{kHz}$, though their amplitude steadily diminishes with increasing frequency, a behaviour indicative of overtones of the epicyclic frequency.
@@
Figure 4: Characteristic GW spectra for models SR (red), IR (black), and FR (yellow), assuming a source distance of $1\,\mathrm{Mpc}$, compared with the design sensitivity curves of current and next-generation interferometers. $h_{\texttt{char}}$ has been convolved over a window of $10\,\mathrm{Hz}$ for better visualization.
@@
$h_{\textnormal{char}}(f)=2f\,\lvert\tilde{h}(f)\rvert$, (2)
where $\tilde{h}(f)$ is the Fourier transform of the GW strain $h(t)$ after applying a Hann window $w(t)$,
$\tilde{h}(f)=\int_{-\infty}^{+\infty}h(t)\,w(t)\,e^{-2\pi ift}\,\mathrm{d}t.$ (3)
The Hann window is defined as
$w(t)=\frac{1}{2}\left[1-\cos\!\left(\frac{2\pi t}{T}\right)\right],\qquad 0\leq t\leq T,$ (4)
and $w(t)=0$ outside the interval $[0,T]$, where $T$ is the total duration of the signal.
The use of a Hann window reduces spectral leakage associated with the finite time span of the signal, but it also introduces a reduction of the Fourier amplitudes. For a coherent signal, the Hann window has a coherent gain $\langle w\rangle=T^{-1}\!\int_{0}^{T}w(t)\,\mathrm{d}t=1/2$ [49]. To preserve the correct normalization of the Fourier amplitude, and hence of $h_{\textnormal{char}}$, the windowed Fourier transform is therefore renormalized by the inverse of this coherent gain.
In Figure 4 we show $h_{\textnormal{char}}/f^{1/2}$ for models SR, IR, and FR at $\mathcal{D}=1\,\mathrm{Mpc}$, juxtaposed with the amplitude spectral density (ASD) of current and next-generation interferometers, namely Advanced LIGO [11, 50, 51], Advanced Virgo [12, 50, 51], KAGRA [13, 50, 51], ET [52, 53], and CE [54, 55].
@@
Model IR differs significantly from the others. It displays two low-frequency spectral humps: one at intermediate frequencies $(60-200\,\mathrm{Hz})$, and another below $30\,\mathrm{Hz}$. The most striking feature of IR is the unmistakable resonance imprint on its characteristic strain. This model yields the strongest signal, with a peak strain more than twice that observed in SR, translating into a peak ASD more than 10 (100) times larger than in model SR (FR)-(see Figure 4). Four prominent peaks, spanning from $\sim 600\,\mathrm{Hz}$ to $\sim 1\,\mathrm{kHz}$, mark the resonant frequencies during the first and second resonant phases (see Figure 1). Overtones at $f\gtrsim 2\,\mathrm{kHz}$ and $f\gtrsim 3\,\mathrm{kHz}$ are also present. To estimate the maximum distance at which the GWs from our models remain detectable, we adopt an optimal signal-to-noise ratio (SNR) threshold of 12. This value is higher than the usual threshold of 8 used in recent works [30, 56] to account for the intrinsic overestimation resulting from the axisymmetric nature of the simulation. We define the SNR as in Moore et al. [48],
$\textnormal{SNR}=\sqrt{\int_{0}^{+\infty}\textnormal{d}\ln f\left(\frac{h_{\textnormal{char}}(f)}{h_{n}(f)}\right)^{2}},$ (5)
where $h_{n}(f)=\sqrt{f}\textnormal{ASD}(f)$, in which we used the ASD of the detector noise. The strong resonant features of model IR lie within the detection range of current interferometers such as Advanced LIGO and Virgo beyond the Andromeda Galaxy [57], with a maximum reach of $1\,\mathrm{Mpc}$. Next-generation observatories like CE, with enhanced high-frequency sensitivity, could extend this range to $\sim 7.4\,\mathrm{Mpc}$.
@@
where $S_{j}(f,t)$ ($j=\textnormal{GW},\,\nu_{i}$) denotes the magnitude of the spectrogram computed with a $10\,\mathrm{ms}$ time window, and $\langle\cdot|\cdot\rangle$ represents the inner product in frequency space.